Bridgeport Inc. has $652,920 to invest. The company is trying to decide between two alternative uses of the funds. One alternative provides $97,304 at the end of each year for 10 years, and the other is to receive a single lump-sum payment of $2,027,870 at the end of the 10 years. Which alternative should Bridgeport select? Assume the interest rate is constant over the entire investment.
Answer #1

Alternative which have higher IRR will be selected.

IRR of first alternative = 652920/97304 = 6.710

IRR = 8%

IRR of second alternative = 652920/2027870 = 0.322

IRR = 12%

Second alternative should be selected.